During 2016, Sarasota Corporation spent $164,160 in research and development costs. As a result, a new product called the New Age Piano was patented. The patent was obtained on October 1, 2016, and had a legal life of 20 years and a useful life of 10 years. Legal costs of $36,720 related to the patent were incurred as of October 1, 2016.

On June 1, 2018, Sarasota spent $9,720 to successfully prosecute a patent infringement suit. As a result, the estimate of useful life was extended to 12 years from June 1, 2018. Prepare all journal entries required in 2018 and 2019.
Answer #1
Date Account Titles and Explanation Debit Credit
Jun. 1, 2018 Patents 9720
Cash 9720
(To record costs incurred on patent infringement suit)
Dec. 31, 2018 Amortization expense 3490
Patents 3490
(To record amortization expense)
Dec. 31, 2019 Amortization expense 3360
Patents 3360
(To record amortization expense)

Working:

Annual amortization on legal costs incurred = $36720/10 = $3672
Amortization expense from Oct. 1, 2016 to Dec. 1, 2016 = $3672 x 3/12 = $918
Amortization expense for 2017 = $3672
Amortization expense from Jan. 1, 2018 to May 31, 2018 = $3672 x 5/12 = $1530
Carrying value of Patents on June 1, 2018 = ($36720 - $918 - $3672 - $1530) + $9720 = $40320
The carrying value will be amortized over the revised estimated useful life of 12 years.
Annual amortization = $40320/12= $3360
Amortization expense from June 1, 2018 to Dec. 31, 2018 = $3360 x 7/12 = $1960
Amortization expense for 2018 = $1530 + $1960 = $3490